Find roots of x
e^x+x=4
1
Expert's answer
2020-10-01T12:22:21-0400

"e^x +x = 4"


Now we write this in a Lamber form i.e.


"1 = (4-x) . e^{-x}"


Rewrite this introducing u i.e.

let "u = 4 - x" such that;

"x = -u +4"

Replacing this back into the Lambert form, we have;


"1 = (4 - (-u+4)) . e^{-(-u+4)}"


"1= (\cancel{4} + u - \cancel{4}) .e^{u-4}"


"1 = u.e^{u-4}" Rewrite this again in lambert form;


"u.e^u = e^4" Solving this, we have;


"u.e^u = e^4 \u21d2 u = W_n (e^4)"


Where W is the Lambert function called the Omega constant


Now substituting this back to "u = - x +4" and solving for "x" ;


"\u21d2 -x +4 = W_n (e^4)"


"\therefore x = - W_n (e^4) +4," "n\isin \Z"